# Security flaws from time-varying active encoding in high-speed measurement-device-independent quantum key distribution

### v1

@@ Abstract
To further understand our key rate results, we can compute the corresponding $\epsilon^{i}_{x}$ values for each case (see Table 1). We can see that the experimentally collected and simulated phase profiles lead to $\epsilon^{i}_{x}$ values on the order of $10^{-3}$. Comparatively, the square phase profiles lead to $\epsilon^{i}_{x}$ values of 0, as they correspond to qubit states. Referring back to Currás-Lorenzo et al. (2023), we can qualitatively understand the impact of $\epsilon^{i}_{x}$ of various orders of magnitude on the secret key rate. For a BB84 protocol using single photons, Currás-Lorenzo et al. (2023) reports a reduction in the maximum secure transmission distance of over 70% (relative to the qubit case) for an $\epsilon^{i}_{x}$ on the order of $10^{-3}$. In addition to our key rate results, this further emphasizes the substantial impact which this side channel has on security.
@@
Naturally, we would expect the time-variance of the phase profiles to have less of an impact on the secret key rate for smaller optical intensity pulse widths. The smaller the optical intensity pulse width, the more constant the temporal phase profile would be over the duration of the pulse width. To confirm this, we computed the secret key rate assuming optical pulse widths (FWHM) of 25 ps and 12.5 ps. The results can be seen in Figure 5. The corresponding $\epsilon^{i}_{x}$ values are also provided in Table 2. As expected, we see dramatic improvements in the secret key rate as the optical pulse width gets smaller in relation to the 200 ps modulation pulse width (along with order of magnitude changes in $\epsilon^{i}_{x}$). This suggests that this side channel can be mitigated by either shortening the optical pulse width or increasing the modulation pulse width. Increasing the modulation pulse width would shorten the duty cycle of the modulation pulse train - this would exacerbate the issue of pulse correlations, which are a major security concern in high clock rate QKD systems Pereira et al. (2020). Shortening the optical pulse width is a relatively viable option - however, for substantial improvements to the secret key rate, techniques such as mode-locking are required for generating pulses on the order of ps Martinez and Yamashita (2011); Zhao et al. (2021).
Figure 5: Key rate results (with respect to the distance between Alice/Bob and the central relay) using the experimentally collected phase profiles (from Figure 3). The results from using the simulated phase profile (from Figure 1) are shown using a dotted line. Here, the width (FWHM) of the Gaussian optical intensity profile is varied. Blue corresponds to a 50 ps width (illustrated in green in Figure 3). Orange corresponds to a 25 ps width and green corresponds a 12.5 ps width. Red corresponds to the case of ideal, time-invariant encoding.
Table 2: Table of $\epsilon^{i}_{x}$ values corresponding to the cases in Figure 5 (varying optical pulse widths).
@@
Conclusion.— In this work, we have identified a side channel which arises in GHz clock rate QKD sources which make use of bandwidth-limited modulation for encoding. At such high clock rates, bandwidth limitations of source equipment start to present several security issues. One such issue is encoding which varies throughout the duration of the optical pulse. This temporal degree of freedom causes the transmitted states to break the qubit assumption, and hence acts as a source of information leakage. This is a side channel which has not been studied in the literature.
We conducted an investigation into this newly identified side channel, using BB84 MDI QKD with decoy states as an illustrative example. Using experimental and simulated results, we demonstrate that this side channel can bring down the secret key rate by several orders of magnitude, as shown in Figure 4.
Although we identified a potential mitigation method, the ideal option would be to remove the time-variance of the phase profiles entirely. At GHz repetition rates, this would be practically unfeasible when using active modulation due to bandwidth limitations. Hence, this work provides further motivation for the use of passive QKD sources, as they circumvent the need for active modulation components Wang et al. (2023); Zapatero et al. (2023); Hu et al. (2023).
Crucially, our work further demonstrates how seemingly minor practical factors in QKD implementations can lead to substantial security risks, particularly when using active modulation components. This highlights the need for experimentalists to be highly aware of imperfections in QKD experiments.
@@ Acknowledgements.
Supplemental Material - Key Rate Analysis Technique
Here, we provide a high level picture for the key rate simulation technique S_Primaatmaja2019 which we employ. This technique bounds the secret key rate by bounding the single photon phase error rate ($e_{ph}$) of the QKD protocol using numerical optimization. The optimization problem can be written as follows.
@@
  • |$\ket{\psi^{i}_{x}}_{A}$ and |$\ket{\phi^{j}_{y}}_{B}$ represent the single photon component of Alice and Bob’s transmitted coherent states, respectively. Superscripts indicate the basis choice (0 denoting the key generation basis and 1 denoting the conjugate basis) while subscripts indicate the encoded bit (0 or 1). Hence, $\braket{\psi^{i^{\prime}}_{x^{\prime}}\varphi^{j^{\prime}}_{y^{\prime}}}{\psi^{i}_{x}\varphi^{j}_{y}}_{AB}$ represent the elements of Alice and Bob’s Gramian matrix S_Primaatmaja2019 . It is through these inner products that the temporal phase profiles enter the key rate analysis.
@@
  • From Equation S1, it is apparent that $\braket{e^{i,j}_{x,y,P}}{e^{i,j}_{x,y,P}}_{E}$ refers to the passing probability of the transmitted state |$\ket{\psi^{i}_{x}\varphi^{j}_{y}}_{AB}$, since |$\ket{e^{i,j}_{x,y,z}}_{E}$ are subnormalized. The variables $p_{\texttt{pass},L}^{i,j,x,y}$ and $p_{\texttt{pass},U}^{i,j,x,y}$ refer to the upper and lower bound on this passing probability. They can be found through the passing probability of Alice and Bob’s overall transmitted coherent state S_Bourassa2022 .
  • $e_{ph}$ can be expressed as follows using the elements of Eve’s Gramian matrix S_Bourassa2022 .
@@
After bounding $e_{ph}$, we can evaluate the secret key rate, $R$. The formula for the asymptotic secret key rate in decoy state MDI QKD is as follows (neglecting the effects of sifting and assuming perfect error correction efficiency) S_Ma2012_mdi_decoy .

### v2

@@ Abstract
To further understand our key rate results, we can compute the corresponding $\epsilon^{i}_{x}$ values for each case (see Table 1). We can see that the experimentally collected and simulated phase profiles lead to $\epsilon^{i}_{x}$ values on the order of $10^{-3}$. Referring back to Currás-Lorenzo et al. (2023), we can qualitatively understand the impact of $\epsilon^{i}_{x}$ of various orders of magnitude on the secret key rate. For a BB84 protocol using single photons, Currás-Lorenzo et al. (2023) reports a reduction in the maximum secure transmission distance of over 70% (relative to the qubit case) for an $\epsilon^{i}_{x}$ on the order of $10^{-3}$. In addition to our key rate results, this further emphasizes the substantial impact which this side channel has on security.
@@
Notice that the smaller the pulse width, the more constant the temporal phase profile would be over the pulse duration, leading to the side channel having a smaller impact. This is confirmed by Figure 4, showing the simulated secret key rates for optical pulse widths (FWHM) of 25 ps and 12.5 ps.
Figure 4: Key rate results (with respect to the distance between Alice/Bob and the central relay) using the experimentally collected phase profiles. The results from using the simulated phase profile are shown using a dotted line. Here, the FWHM of the Gaussian optical intensity profile is varied. Blue: 50 ps width. Orange: 25 ps width. Green: 12.5 ps width. Red: Square phase profiles with ideal amplitudes.
The corresponding $\epsilon^{i}_{x}$ values are also provided in Table 2. As expected, we see dramatic improvements in the secret key rate as the optical pulse width gets smaller in relation to the 200 ps modulation pulse width (along with order of magnitude changes in $\epsilon^{i}_{x}$). This suggests that this side channel can be mitigated by either shortening the optical pulse width or increasing the modulation pulse width. Increasing the modulation pulse width would increase the duty cycle of the modulation pulse train - this would exacerbate the issue of pulse correlations, which are a major security concern in high clock rate QKD systems Pereira et al. (2020). Shortening the optical pulse width is a relatively viable option – however, for substantial improvements to the secret key rate, techniques such as mode-locking are required for generating pulses on the order of ps Martinez and Yamashita (2011); Zhao et al. (2021).
Table 2: Table of $\epsilon^{i}_{x}$ values corresponding to the experimental cases in Figure 4 (varying optical pulse widths).
@@
Conclusion.— In this work, we have identified a new class of side channels which affect quantum cryptographic protocols – in practical experiments, the state preparation process may be entangled with hidden degrees of freedom. We show that this would violate the qubit assumption, opening up a USD attack and severely limiting the maximum secure transmission distance. As an illustrative example, we consider time-varying polarization encoding in GHz clock rate QKD sources exploying active modulation. Through this example, our experimental and simulated side channel characterization revealed that this side channel can reduce the secret key rate by several orders of magnitude, as shown in Figure 3. Although we identified a potential mitigation method for our illustrative example, the optimal solution would be to eliminate the time degree of freedom from the encoding process entirely, ensuring the qubit assumption is not violated. At GHz repetition rates, this would be practically unfeasible when using active modulation due to bandwidth limitations. Hence, this work provides further motivation for the use of passive QKD sources Wang et al. (2023); Zapatero et al. (2023b); Hu et al. (2023). Crucially, our work demonstrates how seemingly minor practical factors in experimentally realized quantum protocols can lead to substantial security risks. This highlights the need for experimentalists to be highly aware of imperfections in experiments.
@@ Acknowledgements.
Supplemental Material
I Key Rate Analysis Technique
A symmetric polarization encoding (BB84) decoy state MDI QKD protocol with intensity settings of 0.05, 0.1, and 0.6 (near typical values Ma et al. (2012)) was assumed in the key rate analysis. Without loss of generality, we consider a projection onto the |$\ket{\Psi^{+}}$ Bell state as a passing measurement. A fiber loss of 0.2 dB/km and a dark count rate of $10^{-6}$/pulse was assumed. We assume that the temporal encoding profiles of Bob’s states are the same as the temporal encoding profiles of Alice’s states.
Now, we provide a high level picture for the key rate simulation technique Primaatmaja et al. (2019) which we employ. This technique bounds the secret key rate by bounding the single photon phase error rate ($e_{ph}$) of the QKD protocol using numerical optimization. The optimization problem can be written as follows.
@@ I Key Rate Analysis Technique
  • |$\ket{\psi^{i}_{x}}_{A}$ and |$\ket{\phi^{j}_{y}}_{B}$ represent the single photon component of Alice and Bob’s transmitted coherent states, respectively. Superscripts indicate the basis choice (0 denoting the key generation basis and 1 denoting the conjugate basis) while subscripts indicate the encoded bit (0 or 1). Hence, $\braket{\psi^{i^{\prime}}_{x^{\prime}}\varphi^{j^{\prime}}_{y^{\prime}}}{\psi^{i}_{x}\varphi^{j}_{y}}_{AB}$ represent the elements of Alice and Bob’s Gramian matrix Primaatmaja et al. (2019). It is through these inner products that the temporal phase profiles enter the key rate analysis. Note that |$\ket{\psi^{i}_{x}}_{A}$ and |$\ket{\phi^{j}_{y}}_{B}$ take the form given by Equation 6 in the main paper.
@@
  • From Equation S1, it is apparent that $\braket{e^{i,j}_{x,y,P}}{e^{i,j}_{x,y,P}}_{E}$ refers to the passing probability of the transmitted state |$\ket{\psi^{i}_{x}\varphi^{j}_{y}}_{AB}$, since |$\ket{e^{i,j}_{x,y,z}}_{E}$ are subnormalized. The variables $p_{\texttt{pass},L}^{i,j,x,y}$ and $p_{\texttt{pass},U}^{i,j,x,y}$ refer to the upper and lower bound on this passing probability. They can be found through the passing probability of Alice and Bob’s overall transmitted coherent state Bourassa et al. (2022).
  • $e_{ph}$ can be expressed as follows using the elements of Eve’s Gramian matrix Bourassa et al. (2022).
@@
After bounding $e_{ph}$, we can evaluate the secret key rate, $R$. The formula for the asymptotic secret key rate in decoy state MDI QKD is as follows (neglecting the effects of sifting and assuming perfect error correction efficiency) Ma et al. (2012).
@@
II Experimental and Simulated Characterization of Temporal Phase Profiles
Here, we describe how we practically acquire the time-varying phase profiles, $\phi^{i}_{x}(t)$, from Equation 6 in the main paper. We acquire the phase profiles using two methods – experiment and simulation. The corresponding key rate results are discussed in the Practical Temporal Phase Profiles section in the main paper.
We start by reviewing electro-optic phase modulation Gee et al. (1983). An electro-optic phase modulator works through a traveling wave mechanism in which an optical signal and traveling RF (radiofrequency) signal propagate together through the modulator with their speeds matched. The voltage required to create a $\pi$ phase change is referred to as $V_{\pi}$ (more generally, we will use $V_{\phi}$ to refer to the voltage required to create a $\phi$ phase change).
As we consider BB84 encoding, we are interested in the temporal phase profiles resulting from the following phase modulations - $\pi/2$, $\pi$, and $-\pi/2$. The width of the phase modulation window (i.e. the width of the voltage pulse used to modulate phase) must be adequately larger than the width of the optical pulse being modulated. Here, we consider a 50 ps wide optical pulse and a 200 ps wide phase modulation window, typically used at GHz clock rates Lu et al. (2021); Kang et al. (2023); Roberts et al. (2018); ichiro Yoshino et al. (2018); Li et al. (2019). Ideally, the voltage pulse used to modulate phase should be square, resulting in time-invariant modulation over the duration of the optical pulse. However, the bandwidth limitations of the arbitrary waveform generator (AWG) and RF amplifier (which is often necessary to create voltage pulses with the required amplitude) result in distorted voltage pulses.
While limited bandwidth is a notable imperfection impacting the time-variance of phase modulation, other equipment-specific factors, such as nonlinearity and the frequency response function, may also contribute. This motivates us to provide optimistically simulated phase profiles in addition to our measured phase profiles as a generalized baseline. The simulated phase profiles solely consider the bandwidth limitations of the AWG, RF amplifier, and electro-optic modulator. We optimistically approximate the frequency response of each component using a simple Butterworth filter of order 2, with -3 dB points of 25 GHz, 11 GHz, and 15 GHz, respectively (corresponding to the equipment specifications). Butterworth filters are an optimistic choice, as their frequency response in the passband is relatively flat. Furthermore, at an order of 2, the frequency response falls off relatively gradually past the passband (in comparison to the frequency response of our state-of-the-art equipment) key (2017); drv (2023); jds (2004). These filters were created and applied to a 200 ps wide square pulse in order to generate the simulated profile (using the SciPy signal processing class) Contributors (2022). The result can be seen in Figure S1.
Figure S1: Simulated $\pi$ phase profile (the simulated ±$\pm\pi/2$ phase profiles are simply a scaled version of this). This was generated by applying Butterworth filters to a 200 ps wide square pulse. The Butterworth filters represent the AWG, amplifier, and electro-optic modulator. The optical pulse intensity profile assumed in this work (50 ps FWHM Gaussian pulse) is also shown in green.
We now move to our experimentally acquired temporal phase profiles. The experimental setup for generating and measuring the phase profiles is shown in Figure S2. The 25 GHz AWG was set to generate 200 ps wide (FWHM) square voltage pulses (the pulse repetition frequency was set to 10 MHz to reduce pulse correlations down to negligible levels). This pulse train was then amplified by an 11 GHz amplifier so that the voltage pulses have the necessary amplitude to generate the desired phase change. Using this setup, voltage pulses with an amplitude of $V_{\pi}$, $V_{\pi/2}$, and $V_{-\pi/2}$ were generated in order to acquire the $\pi$, $\pi/2$, and $-\pi/2$ temporal phase profiles. For our electro-optic modulator, $V_{\pi}$, $V_{\pi/2}$, and $V_{-\pi/2}$ are approximately 5.5 V, 2.75 V, and -2.75 V, respectively.
Figure S2: Experimental setup for generating and measuring phase profiles. The 25 GHz Keysight M8195A AWG and 11 GHz ixBlue DR-VE-10-MO RF amplifier are used to generated voltage pulses which are sent to the JDSU OC-192 intensity modulator (IM). This intensity modulator is a phase modulator with a built in Mach-Zehnder interferometer to convert phase changes to intensity changes. The optical intensity profile (related to the optical phase profile through Equation S4) is collected using the Newport D-8ir 8 ps photodiode (PD) and the 12 GHz Agilent DSO81204A oscilloscope (OSC) after being amplified by a Thorlabs semiconductor optical amplifier (OA). The OA and PD were operated within their linear response region. Yellow indicates an optical fiber connection.
The $\pi$, $\pi/2$, and $-\pi/2$ temporal phase profiles were each acquired by using the corresponding amplified voltage pulses to drive an electro-optic intensity modulator. An electro-optic intensity modulator uses electro-optic phase modulation in conjunction with Mach-Zehnder interferometry in order to modulate intensity. The phase change ($\phi$) applied by the intensity modulator is related to the normalized output intensity ($I$) through Equation S4 when the intensity modulator is biased at a minimum. This relationship can be used to determine a phase profile from its corresponding intensity profile.
$\phi=2cos^{-1}(\sqrt{1-I})$ (S4)
Figure S3: a),b) Voltage pulses used for polarization/phase modulation as well as the corresponding phase profiles. Cubic spline interpolations Contributors (2022) of the profiles are shown using a solid line. The optical pulse intensity profile assumed in this work (50 ps FWHM Gaussian pulse) is also shown in green. Each profile was measured 10 times over an extended time period of several minutes (each time, oscilloscope averaging was used and set to 1024 to average out random noise, such as electronic noise) - the error bars represent the standard deviation in the 10 trials.
The resulting temporal phase profiles are shown in Figure S3 (excluding the $-\pi$/2 profile, which we found to be nearly identical to the $\pi/2$ profile in shape). The corresponding voltage profiles are also shown. Evidently, the $\pi$ and $\pi/2$ phase profiles have a significant level of temporal variation (especially in comparison to the optimistically simulated phase profile from Figure S1, as expected).